how to cancel fractions 3*10*14 /7*3*15
Answer by Alan3354(69443)

3*10*14 /7*3*15
Cancel the 3's
= 10*14/7*15
Factor
= 2*5 * 2*7/7 * 3*5
Cancel the 5 and the 7
= 2*2/3
= 4/3